# Procountable groups are not classifiable by countable structures

### v1

@@ 4. Borel reduction of uniform homeomorphism to isomorphism of procountable groups
Theorem 4.1.
Let $T$ and $U$ be pruned trees on $\mathbb{N}$. Then $T]\cong_{u}[U]$ if and only if $G_{T}$ is topologically isomorphic to $G_{U}$.
@@ 4.1. Proω categories
We need some preliminaries on isomorphism of procountable groups corresponding to certain isomorphisms of inverse systems of type $\mathbb{N}$. We mainly give them to provide a notational framework; more general results are obtained in [20].
@@
Let $\mathcal{K}$ be a category that has as objects countable structures in the same signature, and as morphisms surjective homomorphisms between. The category $\mathrm{Pro}_{\omega}(\mathcal{K})$ has as objects the inverse systems $(\overline{A},\overline{p})=(A_{n},p_{n})_{n\in\mathbb{N}}$ over $\mathcal{K}$, where $p_{n}\colon A_{n+1}\to A_{n}$. The pre-morphisms
@@
are given as sequences of morphisms $f_{n}:A_{\phi(n)}\to B_{n}$ where $\phi:\mathbb{N}\rightarrow\mathbb{N}$ is increasing, and all the diagrams are commutative. In more detail, for $k>n$, write $p_{k,n}$ for the “interval composition” maps $p_{n}\circ p_{n+1}\circ\ldots\circ p_{k-1}\colon A_{k}\to A_{n}$ and define $q_{k,n}$ in a similar way; we require that the diagrams
@@
are commutative. Composition of pre-morphisms is defined in the obvious way (see [20, pg. 5]) and the identity pre-morphism is $({\mathrm{id}_{\overline{A}}},\mathrm{id}_{\mathbb{N}})$, where $\mathrm{id}_{\overline{A}}=(\mathrm{id}_{A_{n}}:n\in\mathbb{N})$, where of course $\mathrm{id}_{A_{n}}:A_{n}\rightarrow A_{n}$ is the identity map on $A_{n}$.
For each pair of objects $(\overline{A},\overline{p})$ and $(\overline{B},\overline{q})$ we define an equivalence relation $\sim$ between pre-morphisms $(\bar{f},\phi),(\bar{f}^{\prime},\phi^{\prime}):(\overline{A},\overline{p})\to(\overline{B},\overline{q})$. We say that $(\bar{f},\phi)\sim(\bar{f}^{\prime},\phi^{\prime})$ provided for each $n\in\mathbb{N}$ there is $m\in\mathbb{N}$ such that $m\geqslant\phi(n),\phi^{\prime}(n)$ and the following diagram is commutative:
@@
Transitivity of $\sim$ and compatibility with composition of pre-morphisms is proved in [20, pg. 6]. The morphisms of $\mathrm{Pro}_{\omega}(\mathcal{K})$ are now defined as the equivalence classes of pre-morphisms under $\sim$.
@@
Notice that for $(\bar{f},\phi):(\overline{A},\overline{p})\to(\overline{A},\overline{p})$, we have $(\bar{f},\phi)\sim({\mathrm{id}_{\overline{A}}},\mathrm{id}_{\mathbb{N}})$ if and only if for every $n\in\mathbb{N}$ there is $m\geqslant\phi(n),n$ such that the following diagram commutes:
$A_{m}$$A_{\phi(n)}$$A_{n}$$A_{n}$$\scriptstyle{p_{m,\phi(n)}}$$\scriptstyle{p_{m,n}}$$\scriptstyle{f_{n}}$$\scriptstyle{\mathrm{id}_{A_{n}}}$
that is, $f_{n}:A_{\phi(n)}\rightarrow A_{n}$ is simply the map $p_{\phi(n),n}$.
Let $\mathrm{EpiGrp}$ be the category of countable groups with onto homomorphisms.
@@
    The inverse limit operation yields a full functor $P$ from $\mathrm{Pro}_{\mathbb{N}}(\mathrm{EpiGrp})$ to procountable groups with continuous epimorphisms.
@@
In order to show that the functor $P$ is full, suppose that $\gamma\colon G\to H$ is a continuous epimorphism. We will define a pre-morphism $(\bar{f},\phi)\colon(\overline{A},\overline{p})\to(\overline{B},\overline{q})$ with $P(\bar{f},\phi)=\gamma$. Let $a_{n}$ and $b_{n}$ be the kernels of the canonical projections $\pi^{G}_{n}:G\to A_{n}$ and $\pi^{H}_{n}:H\to B_{n}$ respectively. Let $\phi$ be an increasing function such that $a_{\phi(n)}\subseteq\gamma^{-1}(b_{n})$. Such $\phi$ exists because $\gamma^{-1}({b_{n}})$ is an open subgroup of $A_{n}$, and the $a_{n}$} form a basis of neighbourhoods of $e$. (This part of the argument is special to procountable groups.) For $n\in\mathbb{N}$ we define $f_{n}:A_{\phi(n)}\rightarrow B_{n}$ via the diagram
@@
We have to verify commutativity of the following diagram, where $k>n$:
@@
In order to do so it suffices to look at the diagram
@@
and recognize its commutativity.
@@
in $\mathrm{Pro}_{\omega}(\mathrm{EpiGrp})$. Then the composition $(\bar{g},\psi)\circ(\bar{f},\phi)$, which we recall to equal
@@
is such that $(\bar{g},\psi)\circ(\bar{f},\phi)\sim(\mathrm{id}_{\overline{A}},\mathrm{id}_{\mathbb{N}})$. By 4.3, this simply means that $g_{n}\circ f_{\psi(n)}=p_{\phi(\psi(n)),n}$. We argue similarly for the composition $(\bar{f},\phi)\circ(\bar{g},\psi)$. Using this, it is easy to see that the following is an isomorphism of topological groups
@@
is an isomorphism in $\mathrm{Pro}_{\omega}(\mathrm{EpiGrp})$. ∎
4.2. Proof of 4.1
Let the maps $p_{n}\colon T_{n+1}\to T_{n}$ and $q_{n}\colon U_{n+1}\to U_{n}$ be given by the parent relation on the trees $T$ and $U$, respectively. This yields inverse sequences $(V_{n},\widehat{p}_{n})_{n\in\mathbb{N}}$ and $(W_{n},\widehat{q}_{n})_{n\in\mathbb{N}}$ over the category $\mathrm{EpiGrp}$. Let $V_{n}=L(T_{n})$ and $W_{n}=L(U_{n})$.
@@
First suppose that $T]\cong_{u}[U]$ via $\Phi$. By 3.5 there are increasing functions $\phi,\psi$ on $\mathbb{N}$ so that for each $x\in[T]$ and $y\in[U]$ we have
$\Phi(x)\upharpoonright n=r_{n}(x\upharpoonright{\phi(n)})$ and $\Phi^{-1}(y)\upharpoonright n=s_{n}(y\upharpoonright{\psi(n)})$,
for some onto functions $r_{n}\colon T_{\phi(n)}\to U_{n}$ and $s_{n}\colon U_{\psi(n)}\to T_{n}$.
We induce morphisms in $\mathrm{Pro}_{\omega}(\mathrm{EpiGrp})$ by pre-morphisms
@@
where $f_{n}=\widehat{r}_{n}$, $q_{n}=\widehat{s}_{n}$. For each $t\in T_{\phi(\psi(n))}$ and $u\in U_{\psi(\phi(n))}$ we have
@@
So $(V_{n},\widehat{p}_{n})_{n}$ and $(W_{n},\widehat{q}_{n})_{n}$ are isomorphic in the category $\mathrm{Pro}_{\omega}(\mathrm{EpiGrp})$. To see this, recall that the binding maps $\widehat{p}_{n}$ and $\widehat{q}_{n}$ act as “predecessor maps”, since so do the maps $p_{n}$ and $q_{n}$. Hence $G_{T}$ and $G_{U}$ are isomorphic by 4.4.
Conversely, suppose that $G_{T}$ and $G_{U}$ are isomorphic. By 4.4 there are pre-morphisms $(\overline{f},\phi)\colon(V_{n},\widehat{p}_{n})_{n}\to(W_{n},\widehat{q}_{n})_{n}$ and $(\overline{g},\psi)\colon(W_{n},\widehat{q}_{n})_{n}\to(V_{n},\widehat{p}_{n})_{n}$ that induce an isomorphism $(V_{n},\widehat{p}_{n})_{n}\to(W_{n},\widehat{q}_{n})_{n}$ and its inverse, respectively, in $\mathrm{Pro}_{\omega}(\mathrm{EpiGrp})$. By virtue of being pre-morphisms, the following diagrams are commutative for each $k>n$:
@@
Because the pre-morphisms induce inverses of each other, we have
Let $r_{n}\colon\subseteq T_{\phi(n)}\to U_{n}$ be defined by
$r_{n}(t)=u$ if $f_{n}(t)$ is conjugate in $W_{n}$ to $u$.
(Note here that $f_{n}(t)^{2}=e$, so $f_{n}(t)$ is conjugate to a unique generator, or it equals $e$. The second case could prima facie happen, so $r_{n}$ could be a partial map, as indicated in the notation.) Similarly, let $s_{n}\colon\subseteq U_{\psi(n)}\to T_{n}$ be defined by
$s_{n}(u)=t$ if $g_{n}(u)$ is conjugate in $V_{n}$ to $t$.
We claim that $s_{n}\circ r_{\psi(n)}=p_{\phi(\psi(n)),n}$ for each $n$. To verify this, let $t\in T_{\phi(\psi(n))}$. By ($\star$4.2) we have that $g_{n}(f_{\psi(n)}(t))\neq e$. So
$f_{\psi(n)}(t)\sim_{W_{\psi(n)}}r_{\psi(n)}(t)\in U_{\psi(n)}$
where $\sim_{L}$ denotes conjugacy in a group $L$. So by ($\star$4.2) again
$\widehat{p}_{\phi(\psi(n)),n}(t)=g_{n}(f_{\psi(n)}(t))\sim_{V_{n}}g_{n}(r_{\psi(n)}(t))\sim_{V_{n}}s_{n}(r_{\psi(n)}(t)).$
Similarly, $r_{n}\circ s_{\phi(n)}=q_{\psi(\phi(n)),n}$. Since the compositions are onto and total, each map $r_{n}$ and $s_{n}$ is onto. Thus the maps $r_{n}$ and $s_{n}$ are also total (composing a partial after an onto map yields a partial map).
Corresponding to the diagrams in EpiGrp above, we have commutative diagrams of onto maps between sets
@@
Hence $T]\cong_{u}[U]$ by 3.5, as required.

### v2

@@ 3.1. Presentations of procountable groups
Remark 3.2.
Proposition 3.3.
Proof.
In the following, letters $s,t$ etc. range over sequences that have an extension in $G$. Recalling 3.1, we claim that (1) is equivalent to the Borel condition
Remark 3.4.
[…]
@@ 3.2. Proω categories
We provide some preliminaries on the correspondence of isomorphisms of procountable groups and certain isomorphisms of inverse systems of type $\omega$. Our purpose is mainly to lay down a notational framework; more general results are obtained in [21] using a somewhat different language.
@@
Let $\mathcal{K}$ be a category that has as objects countable structures in the same signature, and as morphisms surjective homomorphisms between them. The category $\mathrm{Pro}_{\omega}(\mathcal{K})$ has as objects the inverse systems $(\overline{A},\overline{p})=(A_{n},p_{n})_{n\in\mathbb{N}}$ over $\mathcal{K}$, where $p_{n}\colon A_{n+1}\to A_{n}$. The pre-morphisms
@@
are given as sequences of morphisms $f_{n}:A_{\phi(n)}\to B_{n}$ where $\phi:\mathbb{N}\rightarrow\mathbb{N}$ is increasing, and all the diagrams commute. In more detail, for $k>n$, write $p_{k,n}$ for the “interval composition” maps $p_{n}\circ p_{n+1}\circ\ldots\circ p_{k-1}\colon A_{k}\to A_{n}$ and define $q_{k,n}$ in a similar way; we require that the diagrams
@@
commute. Composition of pre-morphisms is defined in the canonical way (see [21, pg. 5]), and the identity pre-morphism is $({\mathrm{id}_{\overline{A}}},\mathrm{id}_{\mathbb{N}})$, where $\mathrm{id}_{\overline{A}}=\{{\mathrm{id}_{A_{n}}}\}_{n\in\mathbb{N}}$.
For each pair of objects $(\overline{A},\overline{p})$ and $(\overline{B},\overline{q})$ one defines an equivalence relation $\sim$ between pre-morphisms $(\bar{f},\phi),(\bar{f}^{\prime},\phi^{\prime}):(\overline{A},\overline{p})\to(\overline{B},\overline{q})$: let $(\bar{f},\phi)\sim(\bar{f}^{\prime},\phi^{\prime})$ if for each $n\in\mathbb{N}$ there is $m\in\mathbb{N}$ such that $m\geqslant\phi(n),\phi^{\prime}(n)$ and the following diagram commutes:
@@
The transitivity of the equivalence relation $\sim$ and its compatibility with composition of pre-morphisms are proved in [21, pg. 6]. The morphisms of $\mathrm{Pro}_{\omega}(\mathcal{K})$ are defined as the equivalence classes of pre-morphisms under $\sim$.
@@
Notice that for a pre-morphism $(\bar{f},\phi):(\overline{A},\overline{p})\to(\overline{A},\overline{p})$, we have $(\bar{f},\phi)\sim({\mathrm{id}_{\overline{A}}},\mathrm{id}_{\mathbb{N}})$ if and only if for every $n\in\mathbb{N}$ there is $m\geqslant\phi(n),n$ such that the diagram above commutes where $\phi^{\prime}(n)=n$, $B_{n}=A_{n}$ and $f_{n}^{\prime}$ is the identity of $A_{n}$. Since the morphisms of $\mathcal{K}$ are surjections, this simply means that $f_{n}=p_{\phi(n),n}$ for each $n$.
Let $\mathrm{GrpEpi}$ be the category of countable groups with onto homomorphisms.
@@
    The inverse limit operation yields a full functor $P$ from $\mathrm{Pro}_{\omega}(\mathrm{GrpEpi})$ to the category of procountable groups with continuous homomorphisms that have dense range.
@@
In order to show that the functor $P$ is full, suppose that $\gamma\colon G\to H$ is a continuous homomorphism with dense range. We will define a pre-morphism $(\bar{f},\phi)\colon(\overline{A},\overline{p})\to(\overline{B},\overline{q})$ with $P(\bar{f},\phi)=\gamma$. Let $a_{n}$ and $b_{n}$ be the kernels of the canonical projections $\pi^{G}_{n}:G\to A_{n}$ and $\pi^{H}_{n}:H\to B_{n}$ respectively. Let $\phi$ be an increasing function such that $a_{\phi(n)}\subseteq\gamma^{-1}(b_{n})$. Such $\phi$ exists because $\gamma^{-1}({b_{n}})$ is an open subgroup of $A_{n}$, and the $a_{n}$} form a basis of neighbourhoods of $e$. (This part of the argument is special to procountable groups.) For $n\in\mathbb{N}$ we define $f_{n}:A_{\phi(n)}\rightarrow B_{n}$ to be the unique map such that the diagram
@@
commutes. Since $\gamma$ has dense range, $\pi^{H}_{n}\circ\gamma$ is onto, and hence $f_{n}$ is onto. To verify that for each $k>n$ the diagram
@@
commutes, it suffices to consider the diagram
@@
and realise that the upper rectangle and the outer rectangle both commute, noting that $p_{\phi(k),\phi(n)}\circ\pi^{G}_{\phi(k)}=\pi^{G}_{\phi(n)}$ and $q_{k,n}\circ\pi^{H}_{k}=\pi^{H}_{n}$.
@@
in $\mathrm{Pro}_{\omega}(\mathrm{GrpEpi})$. Then the composition $(\bar{g},\psi)\circ(\bar{f},\phi)$, which we recall to equal
@@
is such that $(\bar{g},\psi)\circ(\bar{f},\phi)\sim(\mathrm{id}_{\overline{A}},\mathrm{id}_{\mathbb{N}})$. By 3.6, this simply means that $g_{n}\circ f_{\psi(n)}=p_{\phi(\psi(n)),n}$. We argue similarly for the composition $(\bar{f},\phi)\circ(\bar{g},\psi)$. Using this, it is easy to see that the following is an isomorphism of topological groups
@@
is an isomorphism in $\mathrm{Pro}_{\omega}(\mathrm{GrpEpi})$. ∎
4. Borel reduction of uniform homeomorphism to isomorphism of 2-nilpotent procountable groups
In this section we implement the second step of our proof strategy toward a proof of our Main Result 1.2, as laid out in the introduction. In 4.7 we will exhibit a Borel reduction from uniform homeomorphism on Polish ultrametric spaces to isomorphism of nilpotent-2 procountable groups of exponent $p$, where $p$ is an odd prime.
4.1. The 2-nilpotent groups $L(X)$
Let $N$ be the free nilpotent group of class $2$ and exponent $p$ on free generators $\{x_{i}:i\in\mathbb{N}\}$. For distinct $r,s\in\mathbb{N}$ we write
$x_{r,s}=[x_{r},x_{s}].$
As noted in Mekler [22], the centre $Z(N)$ of $N$ is an elementary abelian $p$-group (and hence an $\mathbb{F}_{p}$ vector space) with basis $x_{r,s}$ for $r<s$. Given an undirected irreflexive graph $A$ with vertex set $\mathbb{N}$, Mekler sets
($\diamond$) $G(A)=N/\langle x_{r,s}\colon\,rAs\rangle.$
The centre $Z=Z(G(A))$ is an elementary abelian $p$-group with basis the $x_{r,s}$ such that $\lnot rAs$. Also $G(A)/Z$ is an abelian group of exponent $p$ freely generated by the $Zx_{i}$ (recall that in a $2$-nilpotent group the commutator subgroup is contained in the center). Mekler obtained a normal form for elements of $G(A)$.
Lemma 4.1.
(i) Every element $c$ of $Z$ can be uniquely written in the form
$\prod_{(r,s)\in E}x_{r,s}^{\beta_{r,s}}$
where $E\subseteq\omega\times\mathbb{N}$ is a finite set of pairs $(r,s)$ with $r<s$ and $\lnot rAs$, and $0<\beta_{r,s}<p$.
(ii) Every element of $G(A)$ can be uniquely written in the form $c\cdot v$ where $c\in Z$, and $v=\prod_{i\in C}x_{i}^{\alpha_{i}}$, for $C\subseteq\mathbb{N}$ finite and $0<\alpha_{i}<p$. (The product $\prod_{i\in C}x_{i}^{\alpha_{i}}$ is interpreted along the indices in ascending order.)
The following fact already noted in [22] is evident.
Fact 4.2.
Let $C,D\subseteq\mathbb{N}$. The following holds in $G(A)$.
$\prod_{r\in C}{x_{r}}^{{\alpha}_{r}},\prod_{s\in D}{x_{s}}^{{\beta}_{s}}]=\prod_{r\in C,\,s\in D,\,r<s,\,\lnot Ars}x_{r,s}^{\alpha_{r}\beta_{s}-\alpha_{s}\beta_{r}}$
We will only use Mekler’s construction for a very simple infinite graph: the one such that each connected component has size 2. (While we will no use it, we note that this graph is “nice” in the sense of Mekler, cf. [22, Definition 1.1].)
Notation 4.3.
For $n\in\mathbb{N}$ write $n^{*}=\lfloor n/2\rfloor$. Let $A=\{\langle r,s\rangle\colon\,r\neq s\land r^{*}=s^{*}\}$. Let $L=G(A)$. For $u\in L$ we write $\underline{u}=uZ(L)$, which is an element of the group $L/Z(L)$.
The idea is that $G(A)$ has the minimum commutativity: any pair of elements that commute nontrivially have to generate the same subgroup as some pair $x_{v},x_{v+1}$ for even $v$, all modulo the centre.
Lemma 4.4.
Suppose that $c,d\in L$ are such that $[c,d]=e$.
Either there is an even $i\in\mathbb{N}$ such that $\langle\underline{c},\underline{d}\rangle=\langle\underline{x}_{i},{\underline{x}}_{i+1}\rangle$, or the rank of $\langle\underline{c},\underline{d}\rangle$ as a subgroup of $L/Z(L)$ is at most $1$.
Proof.
We may assume that $c=\prod_{r\in C}{x_{r}}^{{\alpha}_{r}}$ and $d=\prod_{s\in D}{x_{s}}^{{\alpha}_{s}}$ as in 4.1(ii). If $r\in\mathbb{N}\setminus C$ we let $\alpha_{r}=0$, and if $s\in\mathbb{N}\setminus D$ we let $\beta_{s}=0$. Suppose the first condition fails. We show that $\underline{c}=\underline{d}^{\lambda}$ for some $\lambda\in\mathbb{N}$.
By assumption’ there is no $k\in\mathbb{N}$ such that both $C$ and $D$ are contained in $\{2k,2k+1\}$. Note that neither $C$ nor $D$ can be a singleton: if (say) $C=\{r\}$, then there is $s\in D$ such that $r^{*}\neq s^{*}$. Then $\alpha_{r}\beta_{s}\neq 0$ while $\alpha_{s}\beta_{r}=0$, which contradicts the assumption that $c,d]\in Z(L)$ because of 4.2.
Next we claim that $C=D$. Assume for contradiction that $C$ is not contained in $D$ (the case that $D$ is not contained in $C$ is symmetric). Let $r\in C\setminus D$. Then there is $s\in D$ such that $r^{*}\neq s^{*}$ because $D$ is not a singleton. So $\alpha_{r}\beta_{s}\neq 0$ while $\alpha_{s}\beta_{r}=0$. Thus $C=D$.
Now let $m=\min(C)$, and let $\lambda=\alpha_{m}/\beta_{m}$ as an element of $\mathbb{F}_{p}$. If $k\in C$ and $m^{*}<k^{*}$, then $\alpha_{k}/\beta_{k}=\lambda$. Some such $k$ exists. So if $m$ is even and $m+1\in C$, then we also have $\alpha_{m+1}/\beta_{m+1}=\lambda$.
We conclude that $\alpha_{r}=\beta_{r}\lambda$ holds in $\mathbb{F}_{p}$ for each $r\in C$. This shows that $\underline{c}=\underline{d}^{\lambda}$, and hence that the rank of $\langle\underline{c},\underline{d}\rangle$ is at most $1$. ∎
Definition 4.5.
Given a countably infinite set $X$, let $N(X)$ be the free 2-nilpotent group of exponent $p$ on distinct free generators $a_{v},b_{v}$ ($v\in X$). Let
$L(X)=N(X)/\langle\{[a_{v},b_{v}]\colon v\in X\}\rangle.$
Given a set $X$, for $v\in X$ we will interpret $a_{v}$ and $b_{v}$ in $L(X)$. Any surjection $q\colon X\to Y$ induces a group epimorphism $\widehat{q}\colon L(X)\to L(Y)$ given by mapping the equivalence class of $a_{v}$ to the equivalence class $a_{q(v)}$, and similarly for $b_{v}$ and $b_{q(v)}$.
An arbitrary bijection $h\colon X\to\mathbb{N}$ induces an isomorphism $L(X)\to L$ (for $L$ as in 4.3) by sending $a_{v}$ to the equivalence class of $x_{2h(v)}$ and $b_{v}$ to the equivalence class of $x_{2h(v)+1}$ given by the presentation in ($\diamond$4.1). So we can apply Lemma 4.4 to $L(X)$ as well. This lemma can now be used to turn certain group epimorphism $L(X)\to L(Y)$ into surjections $X\to Y$ in a functorial manner. This will be needed to eventually verify that the reduction for the second step (4.7 below) works.
If $r\colon G\to H$ is an epimorphism of groups, we denote by $\underline{r}$ the induced epimorphism $G/Z(G)\to H/Z(H)$.
Lemma 4.6.
Let $\mathcal{L}$ be the category of groups of the form $L(X)$ with morphisms the epimorphism $g\colon L(X)\to L(Y)$ such that for each $v\in X$, the rank of $\langle\underline{g(a_{v})},\underline{g(b_{v})}\rangle$ as a subgroup of $L(Y)/Z(L(Y))$ equals $2$. A functor $F$ from $\mathcal{L}$ to the category of countable sets with surjections is declared by sending $L(X)$ to $X$, and sending a morphism $g:L(X)\to L(Y)$ to the map $p=F(g)\colon X\to Y$ such that
$\langle\underline{g(a_{v})},\underline{g(b_{v})}\rangle=\langle\underline{a_{p(v)}},\underline{b_{p(v)}}\rangle.$
Furthermore, if $q\colon X\to Y$ is a surjection, then $\widehat{q}$ is a morphism of $\mathcal{L}$, and $F(\widehat{q})=q$.
Proof.
To show that $F$ is well-defined for morphisms of $\mathcal{L}$, it suffices to apply 4.4 to $L=L(Y)$, $c=g(a_{v}),d=g(b_{v})$, noting that $[c,d]=e$. One verifies easily that $F$ is compatible with composition, and that $F(\widehat{q})=q$. ∎
4.2. The reduction for the second step
Definition 4.7.
Given a pruned tree $T$, let $T_{n}=\{s\in T\colon|s|=n\}$ be the $n$-th level of the tree $T$, and let $p_{n}\colon T_{n+1}\to T_{n}$ be the predecessor map. We may assume that $T_{1}$ is infinite because the trees we obtain through the coding in 2.4 have that property. Let
[…]
Theorem 4.8.
Let $T$ and $U$ be pruned trees on $\mathbb{N}$ such that $T_{1}$ and $U_{1}$ are infinite. Then $T]\cong_{u}[U]$ if and only if $G_{T}$ is topologically isomorphic to $G_{U}$.
Proof.
Let the maps $p_{n}\colon T_{n+1}\to T_{n}$ and $q_{n}\colon U_{n+1}\to U_{n}$ be given by the predecessor maps on the trees $T$ and $U$, respectively. Write
$V_{n}=L(T_{n})\qquad W_{n}=L(U_{n}).$
So we have inverse systems $(V_{n},\widehat{p}_{n})_{n\in\mathbb{N}}$ and $(W_{n},\widehat{q}_{n})_{n\in\mathbb{N}}$ over the category $\mathrm{GrpEpi}$ determining $G_{T}$ and $G_{U}$, respectively (see 4.5).
@@ 4.2. The reduction for the second step
First suppose that $T]\cong_{u}[U]$ via a uniform homeomorphism $\Phi$. By 2.5 there are increasing functions $\phi,\psi$ on $\mathbb{N}$ and onto functions
$r_{n}\colon T_{\phi(n)}\to U_{n}$ and $s_{n}\colon U_{\psi(n)}\to T_{n}$
so that for each $x\in[T]$ and $y\in[U]$ we have
$\Phi(x)\upharpoonright n=r_{n}(x\upharpoonright{\phi(n)})$ and $\Phi^{-1}(y)\upharpoonright n=s_{n}(y\upharpoonright{\psi(n)})$.
Let $f_{n}=\widehat{r}_{n}$ and $q_{n}=\widehat{s}_{n}$ as defined in 4.5. The pre-morphisms
@@
induce morphisms in $\mathrm{Pro}_{\omega}(\mathrm{GrpEpi})$. For each $t\in T_{\phi(\psi(n))}$ and $u\in U_{\psi(\phi(n))}$,
@@
So $(V_{n},\widehat{p}_{n})_{n}$ and $(W_{n},\widehat{q}_{n})_{n}$ are isomorphic in the category $\mathrm{Pro}_{\omega}(\mathrm{GrpEpi})$ (to see this, recall that the binding maps $\widehat{p}_{n}$ and $\widehat{q}_{n}$ act as “predecessor maps”, since so do the maps $p_{n}$ and $q_{n}$). Hence $G_{T}$ and $G_{U}$ are isomorphic by 3.7.
Conversely, suppose that $G_{T}$ and $G_{U}$ are isomorphic. By 3.7 there are pre-morphisms $(\overline{f},\phi)\colon(V_{n},\widehat{p}_{n})_{n}\to(W_{n},\widehat{q}_{n})_{n}$ and $(\overline{g},\psi)\colon(W_{n},\widehat{q}_{n})_{n}\to(V_{n},\widehat{p}_{n})_{n}$ that induce an isomorphism $(V_{n},\widehat{p}_{n})_{n}\to(W_{n},\widehat{q}_{n})_{n}$ and its inverse, respectively, in $\mathrm{Pro}_{\omega}(\mathrm{GrpEpi})$. Because the pre-morphisms induce inverses of each other, we have
Using the notation of 4.6, the maps on the left hand sides of the equation are morphisms of the category $\mathcal{L}$. We claim that $g_{n}$ and $f_{n}$ are also morphisms of $\mathcal{L}$ for each $n$. We verify it for $g_{n}$, the case of $f_{n}$ being symmetric. We use the left equation in ($\star$4.2). Write $p$ for $p_{\phi(\psi(n)),n}$. For each $v\in T_{\phi(\psi(n)}$, by definition we have $\widehat{p}(\langle a_{v},b_{v}\rangle)=\langle a_{p(v)},b_{p(v)}\rangle$. Thus the rank of $\langle\underline{a}_{v},\underline{b}_{v}\rangle$ cannot be reduced by applying the homomorphism induced by $f_{\psi(n)}$ on the central quotients. So $f_{\psi(n)}$ is in $\mathcal{L}$ by 4.6, and hence $g_{n}$ is in $\mathcal{L}$ by a similar argument. This shows the claim.
Since $(\overline{f},\phi)$ and $(\overline{g},\psi)$ are pre-morphisms, the following diagrams commute for each $k>n$:
@@
Let $r_{n}\colon T_{\phi(n)}\to U_{n}$ and $s_{n}\colon U_{\psi(n)}\to T_{n}$ be defined by $r_{n}=F(f_{n})$ and $s_{n}=F(g_{n})$ where $F$ is the functor from Lemma 4.6. By that lemma we obtain commutative diagrams of surjections between sets
@@
Let $\Phi\colon[T]\to[U]$ and $\Psi\colon[U]\to[T]$ be defined by
$\Phi(x)=\bigcup_{n}r_{n}(x\upharpoonright\phi(n))\qquad\Psi(y)=\bigcup_{n}s_{n}(y\upharpoonright\psi(n)).$
The $\Phi$ and $\Psi$ are uniformly continuous by 2.5. By ($\star$4.2) and 4.6 we have $p_{\phi(\psi(n)),n}=s_{n}\circ r_{\psi(n)}$ and $q_{\psi(\phi(n)),n}=r_{n}\circ s_{\phi(n)}$ for each $n$. Hence $\Psi\circ\Phi$ is the identity on $[T]$, and $\Phi\circ\Psi$ is the identity on $[U]$. So $T]\cong_{u}[U]$ as required. ∎